# The EU AI Act, Stakeholder Needs, and Explainable AI: Aligning Regulatory Compliance in a Clinical Decision Support System

### v1

@@ 6.2 Bridging the AI Act’s Gap
The AI Act deliberately excludes individual patients’ will, focusing instead on roles relevant to product safety. Interestingly, the XAI perspective does not give the patient any more agency than the AI Act does. Thus, both the AI Act and XAI have roles that act on the behalf of the affected party, which by itself lacks the ability, mandate, or obligation to change the functionality of the AI system. However, XAI emphasizes the needs of affected parties in relation to the user, potentially bridging the gap between end users and affected parties, which the AI Act perspective lacks (highlighted as blue arrow in Figure 2(a)).
Notably, the AI Act does not explicitly define stakeholder roles, instead referring to “relevant stakeholders”, “external stakeholders”, “multi-stakeholder” etc. It is therefore our analysis that defines Deployer, Competent Authority and Provider as stakeholder roles. Regarding the equivalence to the XAI roles of User and Affected parties, the AI Act refers to the former as “staff” (the collective that shall be informed, trained and have responsibilities regarding AI systems according to Art. 17 on quality management systems), “affected workers” (those who will use the CE-marked system in relation to Art. 26 and the impact on individuals’ rights) and “natural persons” (those who have the adequate training for using a CE-marked AI system, Art. 26). The latter role is referred to as “categories of natural persons or groups of persons” (impacted by CE-marked systems under Art. 27). As the AI Act lacks systematic terminology for stakeholder roles, we use Natural person and Natural Person(s) to emphasize human involvement (natural person is legal terminology for human) rather than legal persons (legal terminology for organizations, companies, authorities, etc.).
@@ 6.3 From Regulator to Affected Parties—and back
The AI Act mandates respecting the fundamental rights of affected parties but does not specify technical methods or thresholds. This opens a research trajectory for the XAI community in how to achieve appropriate methods and reasonable cut-off points. As an example, how certain does a recommendation system need to be about key-indicators, such as the LOS, that support the decision which patient is moved from the ward to make room for another patient? How do we calculate that certainty and how is it represented to physicians, hospital management, regulators and affected parties? For this to be relevant in a legal context, the future contributions of the XAI community need to be understood and accepted by the diverse group of actors that can take the stakeholder roles of deployer, developer and regulator (see Figure 2(b)). This aligns with Pavlidis (2024), who emphasizes that XAI must operate in compliance settings to meet regulatory requirements. In our case, it led us to target interpretability, usability and trust as key needs to address.
The AI Act introduces operational activities like a data lifecycle and risk management aspects (e.g., the delegation of responsibilities), without specifying methods. Here the XAI community has an opportunity to delimit what is within its interests to contribute towards, but also new aspects of XAI development and usage for future contributions.
@@ 7 Discussion and future research
From the XAI perspective, the starting point of our analysis is the obligations of the deployer. The persons assigned the oversight or usage of the system need to have the right competence, authority, and support so that they can assess if the output is reasonable given the input and current state of the ICU (Art. 26 of the AI Act, which relates to the XAI term of Interpretability). As described in Section 3.2, the CDSS provides key indicator recommendations to support the clinical decision, which patients can be moved to another ward. While XAI can provide measures for explaining how the recommendation came to be, it does not assess whether it is right or wrong. Operating the ward still requires human judgment to assess which recommendations to follow and when alternative routes are more appropriate (see, e.g., (Schoeffer et al., 2024; Gomez et al., 2025) on Human-AI Collaboration).
@@
The overall ambition with the AI Act is to create trust in AI as a technology. It does so by stating general requirements that need to be addressed concretely, both in relation to the intended purpose of the system but also to the involved actors. Here XAI can be of use. As detailed in Section 5.1, the selection of XAI method must be tailored to the stakeholders and their desiderata. The success of this approach depends on meeting the specific desiderata and satisfaction of each end-user. Engaging relevant stakeholders in the selection of XAI method allows addressing their individual desiderata and an effective implementation of different elements of the AI Act.
Yet, there remains much to do. Since the XAI method choices reflect what happens between physician and patient (novice user and affected party), they need to be percolated upwards to the hospital management (deployer) of the AI system and captured as requirements, consequently fulfilled by the system developer (developer) of the AI system. Finally, the design choices and their implementation need to be understood and accepted by the regulator (see Figure 2(b)). XAI can serve as an instrument to build trust at all levels in relation not only to AI as technology but to a specific AI system – the regulator trusts the AI system to be fit for purpose, the deployer and provider trust the AI system to be fit for purpose and the user and the affected parties trust the AI system to be fit for purpose. A similar reasoning is applicable for the usability of the AI system — it needs to be usable for both the end user and the deployer. We foresee a number of studies to further explore the relationship between the levels but also between XAI as a tool and the legal requirements of the AI Act and how they are further detailed in the upcoming harmonized standards.
Here it is worth pointing to an emerging discussion within the academic community on the risks of introducing AI in the healthcare system, specifically in the relationship between clinician and patient and the liability concerns in relation to perceived mistreatment (see, e.g., (Solaiman and Malik, 2024; Duffourc and Gerke, 2023; Čartolovni et al., 2023)). XAI could both serve as a means to lessen the risk of litigation, but also in itself poses a risk if not implemented properly. This is exemplified by Weller (2022), who states that explanations can lead to unsubstantiated trust to the lay user, and hence have a negative impact on affected parties. If XAI is to realize its potential in relation to the AI Act it is important that it is not applied to “XAI-wash” unsuitable systems (see e.g., Rudin (2019) who criticizes the use of XAI methods that may not be suitable in high stakes decisions which are nonetheless justified using XAI), but for a genuine wish to make AI a suitable technology for human needs.

### v2

@@ 6.2 Bridging the AI Act’s Gap
The AI Act deliberately excludes individual patients’ will, focusing instead on roles relevant to product safety. Interestingly, the XAI perspective does not give the patient any more agency than the AI Act does. Thus, both the AI Act and XAI have roles that act on behalf of the affected party, which by itself lacks the ability, mandate, or obligation to change the functionality of the AI system. However, XAI emphasizes the needs of affected parties in relation to the user, potentially bridging the gap between end users and affected parties, which the AI Act perspective lacks (highlighted as blue arrow in Figure˜2(a)).
Notably, the AI Act does not explicitly define stakeholder roles, instead referring to “relevant stakeholders”, “external stakeholders”, “multi-stakeholder” etc. It is therefore our analysis that defines Deployer, Competent Authority, and Provider as stakeholder roles. Regarding the equivalence to the XAI roles of User and Affected parties, the AI Act refers to the former as “staff” (the collective that shall be informed, trained and have responsibilities regarding AI systems according to Art. 17 on quality management systems), “affected workers” (those who will use the CE-marked system in relation to Art. 26 and the impact on individuals’ rights) and “natural persons” (those who have the adequate training for using a CE-marked AI system, Art. 26). The latter role is referred to as “categories of natural persons or groups of persons” (impacted by CE-marked systems under Art. 27). As the AI Act lacks systematic terminology for stakeholder roles, we use Natural person and Natural Person(s) to emphasize human involvement (natural person is legal terminology for human) rather than legal persons (legal terminology for organizations, companies, authorities, etc.).
@@ 6.3 From Regulator to Affected Parties—and back
The AI Act mandates respecting the fundamental rights of affected parties, but does not specify technical methods or thresholds. This opens a research trajectory for the XAI community on how to achieve appropriate methods and reasonable cut-off points. As an example, how certain does a recommendation system need to be about key-indicators, such as the LOS, that support the decision which patient is moved from the ward to make room for another patient? How do we calculate that certainty, and how is it represented to physicians, hospital management, regulators, and affected parties? For this to be relevant in a legal context, the future contributions of the XAI community need to be understood and accepted by the diverse group of actors that can take the stakeholder roles of deployer, developer and regulator (see Figure˜2(b)). This aligns with pavlidis_unlocking_2024, who emphasizes that XAI must operate in compliance settings to meet regulatory requirements. In our case, it led us to target interpretability, usability and trust as key needs to address.
The AI Act introduces operational activities like a data lifecycle and risk management aspects (e.g., the delegation of responsibilities), without specifying methods. Here, the XAI community has an opportunity to delimit what is within its interests to contribute towards, but also new aspects of XAI development and usage for future contributions.
@@ 7 Discussion and future research
From the XAI perspective, the starting point of our analysis is the obligations of the deployer. The persons assigned the oversight or usage of the system need to have the right competence, authority, and support so that they can assess if the output is reasonable given the input and current state of the ICU (Art. 26 of the AI Act, which relates to the XAI term of Interpretability). As described in Section˜3.2, the CDSS provides key indicator recommendations to support the clinical decision, which patients can be moved to another ward. While XAI can provide measures for explaining how the recommendation came to be, it does not assess whether it is right or wrong. Operating the ward still requires human judgment to assess which recommendations to follow and when alternative routes are more appropriate (see, e.g., (schoeffer_explanations_2024; gomez_human-ai_2025) on Human-AI Collaboration).
@@
The overall ambition of the AI Act is to create trust in AI as a technology. It does so by stating general requirements that need to be addressed concretely, both in relation to the intended purpose of the system but also to the involved actors. Here XAI can be of use. As detailed in Section˜5.1, the selection of XAI method must be tailored to the stakeholders and their desiderata. The success of this approach depends on meeting the specific desiderata and the satisfaction of each end-user. Engaging relevant stakeholders in the selection of XAI method allows addressing their individual desiderata and an effective implementation of different elements of the AI Act.
Yet, there remains much to do. Since the XAI method choices reflect what happens between the physician and the patient (novice user and affected party), they need to be percolated upwards to the hospital management (deployer) of the AI system and captured as requirements, consequently fulfilled by the system developer (developer) of the AI system. Finally, the design choices and their implementation need to be understood and accepted by the regulator (see Figure˜2(b)). XAI can serve as an instrument to build trust at all levels in relation not only to AI as technology but to a specific AI system – the regulator trusts the AI system to be fit for purpose, the deployer and provider trust the AI system to be fit for purpose and the user and the affected parties trust the AI system to be fit for purpose. A similar reasoning is applicable for the usability of the AI system — it needs to be usable for both the end user and the deployer. We foresee a number of studies to further explore the relationship between the levels but also between XAI as a tool and the legal requirements of the AI Act and how they are further detailed in the upcoming harmonized standards.
Here, it is worth pointing to an emerging discussion within the academic community on the risks of introducing AI in the healthcare system, specifically in the relationship between clinician and patient and the liability concerns in relation to perceived mistreatment (see, e.g., (solaiman_regulating_2024; Duffourc2023; Cartolovni2023)). XAI could both serve as a means to lessen the risk of litigation, but also, in itself, poses a risk if not implemented properly. This is exemplified by weller_transparency_2022, who states that explanations can lead to unsubstantiated trust to the lay user, and hence have a negative impact on affected parties. If XAI is to realize its potential in relation to the AI Act it is important that it is not applied to “XAI-wash” unsuitable systems (see e.g., rudin_stop_2019 who criticizes the use of XAI methods that may not be suitable in high-stakes decisions which are nonetheless justified using XAI), but for a genuine wish to make AI a suitable technology for human needs.
@@ 8 Conclusion
Statements and Declarations
Competing Interests
The authors have no relevant financial or non-financial interests to disclose.